OpenStudy (needhelpp101):

5/x = 4/8 A. 2 B. 6 C. 8 D. 10

OpenStudy (needhelpp101):

@But

OpenStudy (mathmale):

Might be easier to see what you're supposed to do here if you'd write the problem in this way:\[\frac{ 5 }{ x }=\frac{ 4 }{ 8 }\]

OpenStudy (mathmale):

\[\frac{ 5 }{ x }=\frac{ 4 }{ 8 }\]

OpenStudy (mathmale):

How would you solve for x?

OpenStudy (needhelpp101):

10